# In-Plane Stress Waves in Layered Media: I. Non-Hermitian Degeneracies and Modal Chirality

### v1

@@ Abstract
We study the band structure and scattering of in-plane coupled longitudinal and shear stress waves in layered media and observe that exceptional points (EP) appear for elastic (lossless) media, when parameterized with real-valued frequency and tangential wave vector component. The occurrence of these EP pairs is not limited to the original stop bands. They could also appear in all mode pass bands, leading to the formation of new stop bands. The scattered energy near these locations is studied along with the associated polarization patterns. The broken phase symmetry is observed inside the frequency bands book-ended by these EP pairs. This is especially manifested by the chirality of the trajectory of the particle velocity, which gets selected by a “direction” of the wave, e.g. the imaginary part of normal component of the wave vector, or the energy flux direction just outside the band. Additionally, EP pairs also appear in the spectrum of the (modified) scattering matrix when mechanical gain is theoretically included to balance the loss in a parity-time symmetric finite structure. These EP pairs lead to amplification of transmission to above 1 and single-sided reflectivity, both phenomena associated with broken phase symmetry, with intriguing potential applications.
@@ 2 Transfer matrix of layered structures for oblique coupled P-SV waves
Here we will study the propagation of coupled in-plane shear (SV) and longitudinal (P) stress waves in finite or infinite (periodic) layered solids. Consider a slab consisting of multiple homogeneous and isotropic layers normal to the $x_{1}$-axis. Each layer, $j$, has thickness, $d^{j}=x_{1}^{j}-x_{1}^{j-1}$, where $x_{1}^{j-1}$ and $x_{1}^{j}$ indicate the boundary coordinates of layer $j$, and is unbounded in the two normal directions. Due to the assumed isotropy of each layer and without loss of generality, waves traveling in $x_{1}x_{2}$ plane are studied here.
@@ 2.1 Oblique propagation of coupled P-SV waves
[…]
@@
with $s_{1}^{Pj}=k_{1}^{Pj}/\omega=\sqrt{\left(s^{Pj}\right)^{2}-s_{2}^{2}}$ and $s_{1}^{Sj}=k_{1}^{Sj}/\omega=\sqrt{\left(s^{Sj}\right)^{2}-s_{2}^{2}}$, as the normal components of the slowness vector (wave vector divided by the angular frequency) in which $s^{Pj}=1/c_{L}^{j}$ and $s^{Sj}=1/c_{T}^{j}$ are the longitudinal and shear slowness in layer $j$, respectively [^1] The wave speeds notations $c_{L}$ and $c_{T}$ are consistent with literature, while slowness $s^{Pj}$ and $s^{Sj}$ are used for uniformity in this work, especially as in the laminates the waves are not pure modes. Slownesses are also less used in the literature., $s_{2}$ is the tangential component of the slowness vector and is proportional to the wave vector component $k_{2}=\omega s_{2}$ also parallel to the layers (or the non-dimensional phase $Q_{2}=k_{2}d$, where $d$ is the total thickness of a repeating unit cell). Both are required to be continuous throughout the system due to the Snell’s law. However, formulating based on $s_{2}$ leads to clearer results, especially as $\omega$ approaches 00222Prescribing finite values of $k_{2}$ (or $Q_{2}$), would lead to the appearance of stop bands at zero frequency as the total size of the wave vector $|\bm{k}|$ should also approach zero. In contrast, when finite $s_{2}$ is prescribed, this issue no longer occurs as $k_{2}$ is proportional to the frequency and also vanishes. Therefore while some examples with prescribed $k_{2}$ (or $Q_{2}$) are shown below, for the analysis and major results, $s_{2}$ and $\omega$ are used as the main parameters. In the following formulation the parameterization is dropped for sake of brevity, unless needed to avoid confusion.. The vector $\bm{A}^{j}$ collects the complex amplitude of left and right traveling P- and SV-waves in layer $j$ (first two components are P-wave amplitudes). The definitions of characteristic impedances are as follows: $Z_{1}^{Pj}=C_{11}^{j}s_{1}^{Pj}$, $Z_{12}^{j}=C_{12}^{j}s_{2}$, $Z_{1}^{Sj}=C_{11}^{j}s_{1}^{Sj}$, $Z_{6}^{Pj}=C_{66}^{j}s_{1}^{Pj}$, $Z_{62}^{j}=C_{66}^{j}s_{2}$, and $Z_{6}^{Sj}=C_{66}^{j}s_{1}^{Sj}$.
@@ 3 Mode shapes and scattering near band structure degeneracies
To study the coupled stress wave propagation further and identify occurrence and behavior of various degeneracies in such systems, two example unit cells are introduced in in Figure (1). The base material constants and geometries for each case are shown in Table 1. The unit system [mm, $\mu$s, mg] for length, time, and mass are used throughout, unless explicitly stated other wise. This leads to the units of [km/s, GPa, g/cm3, MHz, MRayl] units for wave velocity, elastic moduli, density, frequency, and impedance, respectively. For better presentation, frequency units of kHz will be used.
@@ 3.1 Degeneracies in band structure and mode shapes
[…]
Figure 5: The real and imaginary parts of the phase advance, $Q_{1}^{\alpha}=K_{1}^{\alpha}d$ at $s_{2}=k_{2}/\omega=0.15$ considering 1% loss in the shear wave speed for layers 2 and 4 ($c_{T}^{\prime\prime}/c_{T}^{\prime}=0.01$).
All of the four exceptional points studied for the 3-phase, 5-layer unit cell are located at the first and the second stop bands. In contrast, Figure (6) shows the phase advance at $s_{2}=0.04$ for the 2-phase, 3-layer unit cell studied in [39] with exceptional points located in the second and third pass bands (EP3-1 to -4) [^3] A third pair, EP3-5 and -6 also appears in the fourth stop band, which due to its similarity to the previous case is not discussed further.. Unlike the previous case in which imaginary parts of the phase advance were equal for the width of the “collapse bands”, in this case the real parts match while the imaginary parts only do so at the two limits. Additionally, while the signs of the real parts are the same, the signs of the imaginary part are opposite, indicating appearance of evanescent modes with decaying and increasing amplitudes. This is indicated by the $2\pi$ or $4\pi$ differences in the real part of the phase advance occurring between P and S pairs with opposite flux ($\pm$). To the best of our knowledge, such a broken phase symmetry phenomenon has not been observed before, particularly in coupled stress waves in layered media,
Figure 6: The real and imaginary part of phase advance at $s_{2}=0.04$ for the 2-phase, 3-layer structure studied in [39]. Mode coalescence happens at the end points of the bands marked with number (1) to (3), the first two of which occur in the middle of pass bands, and are associated with pairs of opposite flux.
@@ 3.2 Modal chirality and flux
It is also interesting to take a closer look at the modal behavior and investigate the polarizations more closely below, within, and above the intervals between the exceptional points. One can utilize the complex amplitudes shown in Figure (3) to construct a cyclic trajectory with $e^{i(-\bm{k}.\bm{x}+\omega t)}$ phasor form. Figure (7) shows the trajectory of real particle velocity vector, consisting of $v_{1}$ and $v_{2}$, at the center of unit cell, $x_{2}=0$. This is depicted at a number of frequency steps near EP5-1 and EP5-2 for the lossless 3-phase, 5 layer, unit cell at $s_{2}=0.15$, associated with the band structure shown in Figure (2). Outside the “collapse” region, the two modes are elliptically polarized, with axes in the tangential and normal directions (indicating ±$\pm\pi/2$ phase difference between the two components). At the exceptional points, the two trajectories coalesce. Between the two points the elliptical trajectories become oblique, turning essentially linear at the center of this region. Furthermore, at this point, the chirality of the trajectories flip from clockwise to counter-clockwise. Since in this case both exceptional points are located within the stop bands, one expects zero energy flux, which can be confirmed by writing the $x_{1}$ component of the flux averaged over a period:
@@
Figure 7: The particle velocity trajectories at $x_{2}=0$ for modes P+ and S+ (negative imaginary parts) at a series of frequencies around exceptional points, EP5-1 and EP5-2 ((a) to (e) are at frequencies less than EP5-1, (f) to (j) are at frequencies between EP5-1 and EP5-2, with (f) and (j) being at the EPs, and (k) to (o) are at frequencies larger than that of EP5-2) The CW and CCW arrows represent the similar chirality of the two modes. Before and after the colliding region, there is a $\pi/2$ phase difference between $v_{1}$ and $v_{2}$, which results in vertical and horizontal elliptical trajectories. At the exceptional points the trajectories coalesce, though they are still not exactly circular; See Figure(3(a)), (3(c)). Within the “collapse” region the phase difference is no longer $\pi/2$ leading to oblique elliptical trajectories. At the center of this region the ellipses become two lines, after which chirality of the trajectories flip from CW to CCW. The reverse chirality evolution is observed for the P- and S- modes with positive imaginary parts.
@@
Figure 8: The polarization trajectories near EP3-1 and -2 at the “collapse” region in Figure (6). This region, unlike the previous case, is inside two pass bands. Mode degeneracy and collapse at EP3-1 and EP3-2 happens for waves of similar phase advance but opposing flux. Flux is zero at exceptional points and all the points in between with non-zero imaginary parts of the phase advance. For those frequency points in between ((g), (h), and (i)), only modes with negative imaginary parts are shown. Unlike the previous case, chirality does not flip in the middle of this stop band. At the center of the band, while it appears that the trajectories collapse, the eigenvalues are not the same.
@@ 3.3 Scattering
Scattering coefficient amplitudes for a finite slab consisting of a single 3-phase, 5-layer, cell are shown in Figure(9) as functions of frequency for both P- and SV-wave incidence. The varying angles of incidence are indicated by the tangential component of the slowness vector $s_{2}$, with $s_{2}=0.48$ equivalent to about $\pi/3$ incidence angle. Figure (10) shows plots of the absorption for these cases. Resonance peaks are varying with the nature of the waves and angle of incidence. The presence of exceptional points in the band structure do not seem to affect the scattering significantly as the main features are associated with the pass bands.
Figure 9: Reflections and transmission amplitudes off of the 3-phase, 5-layer, medium with one unit cell ($n_{c}=1$) for incident plane: (9(a)) P-wave and (9(b)) SV-wave. Note that 1% loss in both wave speeds is added for layers 2 and 4 as the ratio of the imaginary to real parts of the wave speeds ($c^{\prime\prime}/c^{\prime}=0.01$).
Figure 10: Absorption in a single 3-phase, 5-layer, unit cell: (10(a)) P-wave incidence, (10(b)) SV-wave incidence. Layers 2 and 4 have 1% loss in both wave speeds ($c^{\prime\prime}/c^{\prime}=0.01$).
4 Atypical scattering behavior
We now turn our attention from the topology of the band structure to atypical scattering behavior that can be studied in a similar mathematical framework. The scattering matrix, $\mathsf{S}$, is usually organized to calculate reflection and transmission (outgoing) coefficients at each “port/mode” through its operation on a vector representing the incidence (incoming) complex amplitudes of the same ordered set. Here ports are the two semi-infinite domains $a/b$, while modes are the P- or SV-waves in those domains, $P/S$. Ordering first by mode and then by port the scattering matrix can be written as
@@ 4 Atypical scattering behavior
is the 2-by-2 exchange matrix. In this form $\tilde{\mathsf{S}}$ transforms incoming waves from each domain and modality into outgoing waves of the same modality traveling in the same direction but in the other domain. Such a setup lends itself very well to the study of phenomena such as lasing and asymmetric reflection which can occur when phase symmetry is broken, as has been shown earlier for simpler 2 DOF systems in [9] and elsewhere. The eigenvectors of this operator represent combinations of incoming waves that will get transmitted with the same exact complex ratios (i.e. eigen-polarizations) after interacting with the specimen, while the eigenvalues represent the amplification/reduction factor or the phase advance. Therefore, the eigenvalues for a fully elastic or passive lossy specimen must have amplitude 1 (effectively only a phase advance) or less than 1, respectively. For a system with loss and gain in a $\mathcal{PT}$-symmetric arrangement, generally, the 4 eigenvalues still have amplitude 1, though exceptional points of the $\tilde{\mathsf{S}}$ operator can change this and lead to broken phase symmetry. In what follows, we present examples of the appearance of EP pairs in the spectrum of $\tilde{\mathsf{S}}$ along with the geometry and behavior of the ensuing branches. The scattering coefficients in these regions are calculated to evaluate amplification (which may loosely be referred to as “lasing”), as well as asymmetric reflection.
Figure 11: Location of the 4 eigenvalues of the modified scattering matrix $\tilde{\mathsf{S}}$ for the 3-phase, 5-layer, unit cell ($s_{2}=0.2$ in all cases). (11(a)) Eigenvalues are always located on the unit circle for the lossless case and (11(b)) inside it for the lossy case ($c^{\prime\prime}/c^{\prime}=0.01$ in layers 2 and 4). The case of balanced gain and loss ($c^{\prime\prime}/c^{\prime}=\mp 0.05$ in layers 2/4) is shown in (11(c)) and (11(d)) for better view. Note the broken phase symmetry and the higher than 1 amplitude in the second quadrant.
@@
Figure 12: An example of the effect of varying the wave vector component $s_{2}$ and balanced gain parameter $g$ on the eigenvalues of the modified scattering matrix $\tilde{\mathsf{S}}$. The same unit cell as previous figures is studied here. Each case is presented slightly differently for clarity. (12(a)) through (12(c)) show the significant geometrical changes associated with slight variation in wave vector. The latter appears to show a regular crossing, though this needs further high resolution or more detailed study. It also shows a double-lobed feature suggesting mixing of the EP pairs, or possibly a sign of existence of a quadruple EP in nearby complex domain of the parameter space. Comparison of (12(b)) and (12(d)) underlines the sensitivity to the gain parameter $g$, as a reduction in gain turns two elongated regions of broken symmetry into two.
@@
Additionally, due to the higher dimensionality of this operator compared to those studied before (e.g. in [9]), multiple EP pairs may appear in the vicinity of each other. Since the $\mathcal{PT}$-symmetry necessitates that the eigenvalues of $\tilde{\mathsf{S}}$ to always appear in pairs such that $\sigma\sigma^{\prime}=1$, triple branching is not possible. While we have not observed any quadruple branching, we are nonetheless unable to reject their existence based on theoretical considerations readily. However, it may require perfect parametric conditions, unlike the prevalent double branching which appears to occur quite often and for broad ranges of balanced gain ratio. A more accessible study involves bringing multiple EP pairs into the vicinity of each other by exploring the $s_{2}$ space as well as the gain ratio $g=c^{\prime\prime}/c^{\prime}$, denoted here as a negative value for the layers in which gain occurs (the loss value in the corresponding symmetric layers are positive and equal to $-g$). Such adjustments lead to exaggerated geometries of the eigenvalue traces, such as those shown as examples in Figure 12. Additionally, one observes what appears to be regular crossing, which, given the apparent predominance of either level repulsion or exceptional points, is somewhat surprising. It must be emphasized that at very high magnification the presumed regular crossing may prove to be a very small level repulsion or an EP pair. This may be rigorously studied through a number of methods, such as encircling the crossing in the complex $s_{2}$ space [7] or study of the eigenvectors or eigenvector subspaces [47, 48]. Another interesting observation is the emergence of double-lobed features resembling the transition from one EP pair behavior to another neighboring one, or possibly a sign of existence of quadruple branching in the complex parameter domain. One can hypothesize that with minute adjustments to the geometry and/or material properties (including $g$ value), this quadruple branching may be brought onto the real parameter axis, though such a study is not carried out here. Other esoteric geometries show up as well, though their physical significance and/or potential features for application are unclear at the moment.
@@
Figure 13: An example of scattering around EP pairs. The same unit cell as previous figures is studied here, with $s_{2}=0.1695$, $g=-0.05$. (13(a)) and (13(b)) show the scattering amplitudes for an incident longitudinal wave, while (13(c)) and (13(d)) show them for shear incidence. Note the difference in reflection coefficients in $ab$ vs. $ba$ directions, as it diminishes, depending on the direction at a few hundred Hz below and above 8 kHz. Between these two points of single-sided reflection, the amplitude of transmission can be above 1 for $PP$ and $SS$ transmissions but relatively small for cross-polarization transmissions.

### v2

# In-Plane Linear Stress Waves in Layered Media: I. Non-Hermitian Degeneracies and Modal Chirality

@@ Abstract
We study the band structure and scattering of in-plane coupled longitudinal and shear stress waves in linear layered media and observe that exceptional points (EP) appear for elastic (lossless) media, when parameterized with real-valued frequency and tangential wave vector component. The occurrence of these EP pairs is not limited to the original stop bands. They could also appear in all mode pass bands, leading to the formation of new stop bands. The scattered energy near these locations is studied along with the associated polarization patterns. The broken phase symmetry is observed inside the frequency bands book-ended by these EP pairs. This is especially manifested by the chirality of the trajectory of the particle velocity, which gets selected by a “direction” of the wave, e.g. the imaginary part of normal component of the wave vector, or the energy flux direction just outside the band. Additionally, EP pairs also appear in the spectrum of the (modified) scattering matrix when mechanical gain is theoretically included to balance the loss in a parity-time symmetric finite structure. These EP pairs lead to amplification of transmission to above 1 and single-sided reflectivity, both phenomena associated with broken phase symmetry, with intriguing potential applications.
@@ 2 Transfer matrix of layered structures for oblique coupled P-SV waves
Here we will study the propagation of coupled in-plane shear (SV) and longitudinal (P) stress waves in finite or infinite (periodic) layered linear solids. Consider a slab consisting of multiple homogeneous and isotropic layers normal to the $x_{1}$-axis. Each layer, $j$, has thickness, $d^{j}=x_{1}^{j}-x_{1}^{j-1}$, where $x_{1}^{j-1}$ and $x_{1}^{j}$ indicate the boundary coordinates of layer $j$, and is unbounded in the two normal directions. Due to the assumed isotropy of each layer and without loss of generality, waves traveling in $x_{1}x_{2}$ plane are studied here.
@@ 2.1 Oblique propagation of coupled P-SV waves
$mathbold{\psi}(x_{1},\omega,s_{2})=\begin{pmatrix}v_{1}\\ \tau_{1}\\ v_{2}\\ \tau_{6}\end{pmatrix}=\bm{\zeta}^{j}(s_{2})\bm{\delta}^{j}(x_{1},\omega,s_{2})\bm{A}^{j}e^{-\mathrm{i}\omega s_{2}x_{2}},$ (1)
@@
with $s_{1}^{Pj}=k_{1}^{Pj}/\omega=\sqrt{\left(s^{Pj}\right)^{2}-s_{2}^{2}}$ and $s_{1}^{Sj}=k_{1}^{Sj}/\omega=\sqrt{\left(s^{Sj}\right)^{2}-s_{2}^{2}}$, as the normal components of the slowness vector (wave vector divided by the angular frequency) in which $s^{Pj}=1/c_{L}^{j}$ and $s^{Sj}=1/c_{T}^{j}$ are the longitudinal and shear slowness in layer $j$, respectively [^1] The wave speeds notations $c_{L}$ and $c_{T}$ are consistent with literature, while slowness $s^{Pj}$ and $s^{Sj}$ are used for uniformity in this work, especially as in the laminates the waves are not pure modes. Slownesses are also less used in the literature., $s_{2}$ is the tangential component of the slowness vector and is proportional to the wave vector component $k_{2}=\omega s_{2}$ also parallel to the layers (or the non-dimensional phase $Q_{2}=k_{2}d$, where $d$ is the total thickness of a repeating unit cell). Both are required to be continuous throughout the system due to the Snell’s law. However, formulating based on $s_{2}$ leads to clearer results, especially as $\omega$ approaches 0222Prescribing finite values of $k_{2}$ (or $Q_{2}$), would lead to the appearance of stop bands at zero frequency as the total size of the wave vector $|\bm{k}|$ should also approach zero. In contrast, when finite $s_{2}$ is prescribed, this issue no longer occurs as $k_{2}$ is proportional to the frequency and also vanishes. Therefore while some examples with prescribed $k_{2}$ (or $Q_{2}$) are shown below, for the analysis and major results, $s_{2}$ and $\omega$ are used as the main parameters. In the following formulation the parameterization is dropped for sake of brevity, unless needed to avoid confusion.. The vector $\bm{A}^{j}$ collects the complex amplitude of left and right traveling P- and SV-waves in layer $j$ (first two components are P-wave amplitudes). The definitions of characteristic impedances are as follows: $Z_{1}^{Pj}=C_{11}^{j}s_{1}^{Pj}$, $Z_{12}^{j}=C_{12}^{j}s_{2}$, $Z_{1}^{Sj}=C_{11}^{j}s_{1}^{Sj}$, $Z_{6}^{Pj}=C_{66}^{j}s_{1}^{Pj}$, $Z_{62}^{j}=C_{66}^{j}s_{2}$, and $Z_{6}^{Sj}=C_{66}^{j}s_{1}^{Sj}$.
@@ 3 Mode shapes and scattering near band structure degeneracies
[…]
@@ 3.1 Degeneracies in band structure and mode shapes
Figure 4: The real and imaginary parts of the phase advance at different values of $s_{2}$. $s_{2}$ changes from 0 to 0.1820 at 9 levels shown in (4(a)) to (4(i)). The first longitudinal-dominant stop band remain relatively stationary while the shear-dominant one shifts up as $s_{2}$ increases and the simple crossing (at $s_{2}=0$) turn into mode coalescence regions (around $6\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$). The imaginary parts of the phase advance are equal over the entire regions, while real parts match only at the end point, i.e. the locations of the exceptional points. The Collapse region I, at around $15\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$ for $s_{2}=0.15$, shows up sooner and takes longer to evolve, and in the process flips what appears to be longitudinal- and shear-dominant branches twice.
Figure (5) shows the same for a lossy case (1% added loss to both shear and longitudinal wave speeds in layers 2 and 4). The coalescence in $Q_{1}^{\alpha}$ occurs at four different frequencies below $20\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. Note that while the band structure looks similar in the lossy case, one can no longer observe coalescence of complex eigenvalues, and the imaginary part of the eigenvalues simply cross at a single point (where the real parts are different) rather than collapsing for a finite frequency band.
Figure 5: The real and imaginary parts of the phase advance, $Q_{1}^{\alpha}=K_{1}^{\alpha}d$ at $s_{2}=k_{2}/\omega=0.15$ considering 1% loss in the shear wave speed for layers 2 and 4 ($c_{T}^{\prime\prime}/c_{T}^{\prime}=0.01$). Frequency unit is kHz.
All of the four exceptional points studied for the 3-phase, 5-layer unit cell are located at the first and the second stop bands. In contrast, Figure (6) shows the phase advance at $s_{2}=0.04$ for the 2-phase, 3-layer unit cell studied in [39] with exceptional points located in the second and third pass bands (EP3-1 to -4) [^3] A third pair, EP3-5 and -6 also appears in the fourth stop band, which due to its similarity to the previous case is not discussed further.. Unlike the previous case in which imaginary parts of the phase advance were equal for the width of the “collapse bands”, in this case the real parts match while the imaginary parts only do so at the two limits. Additionally, while the signs of the real parts are the same, the signs of the imaginary part are opposite, indicating appearance of evanescent modes with decaying and increasing amplitudes. This is indicated by the $2\pi$ or $4\pi$ differences in the real part of the phase advance occurring between P and S pairs with opposite flux ($\pm$). To the best of our knowledge, such a broken phase symmetry phenomenon has not been observed before, particularly in coupled stress waves in linear layered media,
Figure 6: The real and imaginary part of phase advance at $s_{2}=0.04$ for the 2-phase, 3-layer structure studied in [39]. Mode coalescence happens at the end points of the bands marked with number (1) to (3), the first two of which occur in the middle of pass bands, and are associated with pairs of opposite flux. Frequency unit is kHz.
@@ 3.2 Modal chirality and flux
It is also interesting to take a closer look at the modal behavior and investigate the polarizations more closely below, within, and above the intervals between the exceptional points. One can utilize the complex amplitudes shown in Figure (3) to construct a cyclic trajectory with $e^{i(-\bm{k}.\bm{x}+\omega t)}$ phasor form. Figure (7) shows the trajectory of real particle velocity vector, consisting of $v_{1}$ and $v_{2}$, at the center of unit cell, $x_{2}=0$. This is depicted at a number of frequency steps near EP5-1 and EP5-2 for the lossless 3-phase, 5 layer, unit cell at $s_{2}=0.15$, associated with the band structure shown in Figure (2). Outside the “collapse” region, the two modes are elliptically polarized, with axes in the tangential and normal directions (indicating $\pm\pi/2$ phase difference between the two components). At the exceptional points, the two trajectories coalesce. Between the two points the elliptical trajectories become oblique, turning essentially linear at the center of this region. Furthermore, at this point, the chirality of the trajectories flip from clockwise to counter-clockwise. Since in this case both exceptional points are located within the stop bands, one expects zero energy flux, which can be confirmed by writing the $x_{1}$ component of the flux averaged over a period:
@@
Figure 7: The particle velocity trajectories at $x_{2}=0$ for modes P+ and S+ (negative imaginary parts) at a series of frequencies around exceptional points, EP5-1 and EP5-2 ((a) to (e) are at frequencies less than EP5-1, (f) to (j) are at frequencies between EP5-1 and EP5-2, with (f) and (j) being at the EPs, and (k) to (o) are at frequencies larger than that of EP5-2) The CW and CCW arrows represent the similar chirality of the two modes. Before and after the colliding region, there is a $\pi/2$ phase difference between $v_{1}$ and $v_{2}$, which results in vertical and horizontal elliptical trajectories. At the exceptional points the trajectories coalesce, though they are still not exactly circular; See Figure(3(a)), (3(c)). Within the “collapse” region the phase difference is no longer $\pi/2$ leading to oblique elliptical trajectories. At the center of this region the ellipses become two lines, after which chirality of the trajectories flip from CW to CCW. The reverse chirality evolution is observed for the P- and S- modes with positive imaginary parts. Quiver plot animations of selected modes (particle velocity vector) is provided in the supplementary materials.
@@
Figure 8: The polarization trajectories near EP3-1 and -2 at the “collapse” region in Figure (6). This region, unlike the previous case, is inside two pass bands. Mode degeneracy and collapse at EP3-1 and EP3-2 happens for waves of similar phase advance but opposing flux. Flux is zero at exceptional points and all the points in between with non-zero imaginary parts of the phase advance. For those frequency points in between ((g), (h), and (i)), only modes with negative imaginary parts are shown. Unlike the previous case, chirality does not flip in the middle of this stop band. At the center of the band, while it appears that the trajectories collapse, the eigenvalues are not the same. Quiver plot animations of selected modes (particle velocity vector) is provided in the supplementary materials.
@@ 3.3 Scattering
Scattering coefficient amplitudes for a finite slab consisting of a single 3-phase, 5-layer, cell are shown in Figure(9) as functions of frequency for both P- and SV-wave incidence. The varying angles of incidence are indicated by the tangential component of the slowness vector $s_{2}$, with $s_{2}=0.48$ equivalent to about $\pi/3$ incidence angle. Figure (10) shows plots of the absorption for these cases. Resonance peaks are varying with the nature of the waves and angle of incidence. The presence of exceptional points in the band structure do not seem to affect the scattering significantly as the main features are associated with the pass bands. The high transmission / low reflection in P-wave spectrum clearly matches with the beginning of the second longitudinal pass band (“optical branch”), which is relatively fixed around $8\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$, and the same can be observed for the second shear pass band as it changes with $s_{2}$; See Figure (4). Furthermore, this also corresponds with increased dissipation inside the system as indicated in Figure (10). Finally for higher oblique angles (higher $s_{2}$), any incident wave would activate both shear and longitudinal waves inside the layered medium, which in turn lead to both longitudinal and shear waves transmission and reflection in the exterior domains, as indicated in the coupled scattering coefficients as seen in Figure (9), with both effects becoming more prominent at more oblique angles.
Figure 9: Reflections and transmission amplitudes off of the 3-phase, 5-layer, medium with one unit cell ($n_{c}=1$) for incident plane: (9(a)) P-wave and (9(b)) SV-wave. Note that 1% loss in both wave speeds is added for layers 2 and 4 as the ratio of the imaginary to real parts of the wave speeds ($c^{\prime\prime}/c^{\prime}=0.01$). Frequency unit is kHz.
Figure 10: Absorption in a single 3-phase, 5-layer, unit cell: (10(a)) P-wave incidence, (10(b)) SV-wave incidence. Layers 2 and 4 have 1% loss in both wave speeds ($c^{\prime\prime}/c^{\prime}=0.01$). Frequency unit is kHz.
4 Scattering behavior in $\mathcal{PT}$-symmetric media
We now turn our attention from the topology of the band structure to the study of the scattering response with a focus on potential for lasing (requiring wave amplification and divergence of mode lifetime [47]) and and highly asymmetric reflection. While the former requires active (gain) domains, the latter can be simply achieved through asymmetric structures, i.e. when $\mathcal{P}$-symmetry is broken. Here we show the occurrence of both wave amplification and asymmetric reflection in $\mathcal{PT}$-symmetric linear media. Such phenomena can occur when phase symmetry is broken, as has been shown earlier for simpler 2 DOF systems in [9] and elsewhere. In the cases studied below, a transmission coefficient with amplitude greater than 1 (through a single cell) is expected through the increase of the (balanced) gain. Furthermore, we show a difference of more than two orders of magnitude in directional reflection coefficients with very small and equal (i.e. balanced) gain and loss factors.
The scattering matrix, $\mathsf{S}$, is usually organized to calculate reflection and transmission (outgoing) coefficients at each “port/mode” through its operation on a vector representing the incidence (incoming) complex amplitudes of the same ordered set. Here ports are the two semi-infinite domains $a/b$, while modes are the P- or SV-waves in those domains, $P/S$. Ordering first by mode and then by port the scattering matrix can be written as
@@ 4 Scattering behavior in $\mathcal{PT}$-symmetric media
is the 2-by-2 exchange matrix. In this form $\tilde{\mathsf{S}}$ transforms a vector consisting of the complex amplitudes of incoming waves from each domain and modality into a vector consisting of the complex amplitudes of outgoing waves of ordered with the same modality traveling in the same direction but in the other domain. Such a rearrangement lends itself very well to the focused study of this section. The eigenvectors of this operator represent combinations of incoming waves that will get transmitted with the same exact complex ratios (i.e. eigen-polarizations) after interacting with the specimen, while the eigenvalues represent the amplification/reduction factor or the phase advance. Therefore, the eigenvalues for a fully elastic or passive lossy specimen must have amplitude 1 (effectively only a phase advance) or less than 1, respectively. For a system with loss and gain in a $\mathcal{PT}$-symmetric arrangement, generally, the 4 eigenvalues still have amplitude 1, though exceptional points of the $\tilde{\mathsf{S}}$ operator can change this and lead to broken phase symmetry. In what follows, we present examples of the appearance of EP pairs in the spectrum of $\tilde{\mathsf{S}}$ along with the geometry and behavior of the ensuing branches. The scattering coefficients in these regions are calculated to evaluate transmission amplification factor above 1 as well as highly asymmetric reflection.
Figure 11: Location of the 4 eigenvalues of the modified scattering matrix $\tilde{\mathsf{S}}$ for the 3-phase, 5-layer, unit cell ($s_{2}=0.2$ in all cases). All subplots (a–d) correspond to the same frequency range, and the color bar shown (in kHz units) applies to all. (11(a)) Eigenvalues are always located on the unit circle for the lossless case and (11(b)) inside it for the lossy case ($c^{\prime\prime}/c^{\prime}=0.01$ in layers 2 and 4). The case of balanced gain and loss ($c^{\prime\prime}/c^{\prime}=\mp 0.05$ in layers 2/4) is shown in (11(c)) and (11(d)) for better view. Note the broken phase symmetry and the higher than 1 amplitude in the second quadrant.
@@
Figure 12: An example of the effect of varying the wave vector component $s_{2}$ and balanced gain parameter $g$ on the eigenvalues of the modified scattering matrix $\tilde{\mathsf{S}}$. Panels (a) and (c) share the same frequency range (in kHz units) and color scale as (b). The same unit cell as previous figures is studied here. Each case is presented slightly differently for clarity. (12(a)) through (12(c)) show the significant geometrical changes associated with slight variation in wave vector. The latter appears to show a regular crossing, though this needs further high resolution or more detailed study. It also shows a double-lobed feature suggesting mixing of the EP pairs, or possibly a sign of existence of a quadruple EP in nearby complex domain of the parameter space. Comparison of (12(b)) and (12(d)) underlines the sensitivity to the gain parameter $g$, as a reduction in gain turns two elongated regions of broken symmetry into two.
@@
Additionally, due to the higher dimensionality of this operator compared to those studied before (e.g. in [9]), multiple EP pairs may appear in the vicinity of each other. Since the $\mathcal{PT}$-symmetry requires the eigenvalues of $\tilde{\mathsf{S}}$ to always appear in pairs such that $\sigma\sigma^{\prime}=1$, triple branching is not possible. Although we have not observed any quadruple branching, we are nonetheless unable to readily reject their existence based on theoretical considerations. However, it may require perfect parametric conditions, unlike the prevalent double branching which appears to occur quite often and for broad ranges of balanced gain ratio. A more accessible study involves bringing multiple EP pairs into the vicinity of each other by exploring the $s_{2}$ space as well as the gain ratio $g=c^{\prime\prime}/c^{\prime}$, denoted here as a negative value for the layers in which gain occurs (the loss value in the corresponding symmetric layers are positive and equal to g-g). Such adjustments lead to exaggerated geometries of the eigenvalue traces, similar to those shown as examples in Figure 12. Additionally, one observes what appears to be regular crossing, which, given the apparent predominance of either level repulsion or exceptional points, is somewhat surprising. It must be emphasized that at very high magnification the presumed regular crossing may prove to be a very small level repulsion or an EP pair. This may be rigorously studied through a number of methods, such as encircling the crossing in the complex $s_{2}$ space [7] or study of the eigenvectors or eigenvector subspaces [48, 49]. Another interesting observation is the emergence of double-lobed features resembling the transition from one EP pair behavior to another neighboring one, or possibly a sign of existence of quadruple branching in the complex parameter domain. One can hypothesize that with minute adjustments to the geometry and/or material properties (including $g$ value), this quadruple branching may be brought onto the real parameter axis, though such a study is not carried out here. Other esoteric geometries show up as well, though their physical significance and/or potential features for application are unclear at the moment.
@@
Figure 13: An example of scattering around EP pairs. The x-axis in all subplots represents frequency (in kHz), and the y-axis shows the magnitude of the scattering coefficients. The same unit cell as previous figures is studied here, with $s_{2}=0.1695$, g=0.05g=-0.05. (13(a)) and (13(b)) show the scattering amplitudes for an incident longitudinal wave, while (13(c)) and (13(d)) show them for shear incidence. Note the difference in reflection coefficients in abab vs. baba directions, as it diminishes, depending on the direction at a few hundred Hz below and above $8\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. Between these two points of single-sided reflection, the amplitude of transmission can be above 1 for PPPP and SSSS transmissions but relatively small for cross-polarization transmissions.